OpenStudy (anonymous):

1. Max's age is 2 more than his father's age divided by 4. Max is 13 years old. How old is his dad? please show the equation

OpenStudy (cwrw238):

let max age = x and fathers = y x =( y/5 ) + 2 x = 13 so 13 = ( y/5 ) + 2 y/5 = 11 y = 55 his father is 55

OpenStudy (anonymous):

Suppose max father age is 5x, and age of max is y: So, \(y = x + 2\) \(y = 13\) (Given) So, \(x = 13 - 2 = 11\) So his father age is : \(5x = 5 \times 11 = 55\)